Shiv want to lay a track 200 m long and 6 m wide using a team of 60 boys in 2 days. After 1 day, Shiv find 75 m x 6 m track has been laid. How many men should Shiv hire additionally if 2 men do the job of 5 boys


1
Expert's answer
2021-03-17T17:12:22-0400

Solution:

Since, width is same all the time, so we are avoiding it and taking only length into account.

60 boys lay 200 m track in 2 days.

60 boys lay 100 m track in 1 day.

1 boy lays "(\frac {100}6=\frac 53)" m track in 1 day.

After 1 day, 75 m track is laid.

Remaining track = 200 - 75 = 125 m

Number of boys required to lay this "=125\times \frac35=75"

"=15\times 5\ boys
\\=15\times 2\ men \ [\because 5\ boys=2 \ men]
\\=30\ men"

So, 30 men should be hired additionally.